Debbie has 8% more money than Marty. Mary has 4% more money than Marty. They all have a total of $377. How much money does Marty have?
Answer by checkley79(3341)

MARTY=X
MARY=Y
DEBBIE=Z
Y=1.02X
Z=1.08X
X+Y+Z=377
X+1.02X+1.08X=377
3.1X=377
X=377/3.1
X=$121.61 FOR MARTY
1.02*121.61=$124.05 FOR MARY
1.08*121.61=$131.34 FOR DEBBIE
PROOF:
121.61+124.05+131.34=377
377=377